How many liters of a 20% alcohol should be added to an 80% alcohol solution to obtain 15 liters of a 60% alcohol solution?
Answer by stanbon(75887)

How many liters of a 20% alcohol should be added to an 80% alcohol solution to obtain 15 liters of a 60% alcohol solution?
------
Equation::
alch + alch = alch
----
0.20x + 0.80(15-x) = 0.60*15
-----
20x + 80*15 - 80x = 60*15
-60x = -20*15
x = (1/3)15
x = 5 liters (amt. of 20% solution needed)
------------
Cheers,
Stan H.
-----------------